OpenStudy (anonymous):

Integrate 1/x^2-4 dx

OpenStudy (anonymous):

\[\int\limits_{}^{}\frac{ 1 }{ x^2-4 }dx\]

OpenStudy (anonymous):

make partial fractions

OpenStudy (anonymous):

i don't know partial fractions

OpenStudy (anonymous):

couldn't i just make it; \[(x^2-4)^{-1}\]

OpenStudy (anonymous):

oh but then i would divide by zero

OpenStudy (anonymous):

oh nevermind i would get: \[\frac{ 1 }{ 0 }\]

OpenStudy (anonymous):

\[\frac{ 1 }{x ^{2}-4 }=\frac{ 1 }{ \left( x-2 \right)\left( x+2 \right) }\] \[=\frac{ A }{x-2 }+\frac{ B }{ x+2 }\] \[1=A \left( x+2 \right)+B \left( x-2 \right)\] equating coefficients of like powers. 0=A+B 1=2A-2B solving for A and B B=-A 1=2A+2A A=1/4 B=-1/4

OpenStudy (anonymous):

ok i kind of get it

OpenStudy (anonymous):

\[\int\limits \frac{ dx }{x ^{2}-4}=\frac{ 1 }{ 4 } \int\limits \frac{ dx }{x-2 }-\frac{ 1 }{ 4 } \int\limits \frac{ dx }{x+2 }\] solve it.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

yw